# Bimanual Regrasp Planning and Control for Active Reduction of Object Pose Uncertainty

### v3

@@ Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concept that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that it had good repeatability. The deviation levels of the experimental trials were on the same order of magnitude as those of an optical tracking system, demonstrating strong relative inference performance.
@@ I Introduction
Figure 1: (a) Precisely grasping an object is difficult. (a.i) Before grasping. (a.ii) After grasping. (a.iii) Comparison. (b) Flat finger pads of a parallel gripper help constrain uncertainty to a plane. (c) Three orthogonal parallel grasps constrain three planes, leading to an exact pose.
This paper proposes a bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to reduce uncertainties. It does not require external fixtures or additional sensors and is more flexible for diverse robotic manipulation scenarios. The fundamental concept of the proposed method is illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state Note 1 Note 1 Note 1This unique pose constraint holds under the following assumptions: i) The object is sufficiently rigid such that the grasping force is transmitted and sensed through the force sensors, rather than causing deformation. ii) The contact between the gripper fingers and the object is assumed to be flat surface contact with sufficient friction to prevent slippage. iii) The maximum uncertainty in the object pose is within the gripper’s opening range, ensuring valid contact during grasping.. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
@@ II-A Categories of Regrasp Planning Methods
Continuous regrasp methods, on the other hand, aim to adjust object poses through controlled rolling or sliding motions while maintaining contact. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds. The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
@@ III Regrasp Planning Using Orthogonal Triplets
Figure 2: (a) Antipodal grasp planning. (b) Candidate grasp poses planned using the antipodal method. (c) Separating the candidate grasp poses into three mutually orthogonal groups considering their opening and closing directions. Each color in the subfigure represents one group.
Subsequently, based on the results of the grasp planning, we separate the grasps into groups. Since our goal is to identify three mutually orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 2 Note 2 Note 2In this context, the direction of $\boldsymbol{v}_{i}$ is not considered.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …. The candidate grasp poses in Fig. 2(b) can be correctly classified into three mutually orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue colors.
After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triplet combination.
@@
Although we consistently refer to “orthogonal” grasps or groups, it is important to note that strict 90-degree orthogonality is not required to determine the object’s pose. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triplet combinations. The triplets may have non-zero scores, However, as long as the grasp direction vectors $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ form a non-singular matrix when arranged as columns, the pose can be estimated. This point will be clearer in Section V when we introduce the pose estimation algorithm. Here, the reason we sort triplet grasp combinations by their score is not for enforcing strict orthogonality, but rather to account for potential changes of grasp directions caused by subsequent admittance control. Prioritizing grasp triplets with higher orthogonality helps minimize the impact of grasp direction deviations. We will discuss this aspect in more detail in the admittance control section.
After identifying orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. We use an incremental method to explore combinations of grasp triplets from the orthogonal groups and output a successful motion once found. Each triplet comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triplets and output successful motion once found. The detailed algorithmic flow is illustrated in Fig. 3.
@@
It should be noted that our grasping and planning methods are probabilistic, and their success rates inherently depend on the granularity, number, and diversity of the grasp samples. At the same time, increasing the number of grasps inevitably impacts planning time. Identifying the optimal trade-off between these factors is an important challenge that could be addressed using predictive approaches. If a large amount of data can be accumulated using the proposed method in the future, we are interested in exploring probabilistic modeling methods to learn grasp triplets and thus improve planning efficiency. However, this is beyond the scope of the present paper and is therefore not elaborated here.
@@ VI Experiments and Analysis
Table I summarizes the results. The data section has two main column groups reporting the positional differences and rotational differences. Within each group, there are three subcolumns. The $g_{2}$ columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control. The $g_{3}$ columns present the corresponding differences for the $g_{3}$ grasp. The “object” columns report the differences between the object pose estimated by the proposed method and the pose measured by the optical tracking system. The results for $g_{1}$ are not included, as it involves no admittance control.
@@
Task Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
$g_{2}$ $g_{3}$ object $g_{2}$ $g_{3}$ object
@@
Upon closer examination, we found that these errors are primarily systematic. For the L-shaped object, the standard deviations in $\Delta\mathbf{p}$ across tasks were [0.25, 0.15, 0.25] mm, all below 0.3 mm, and the deviations in $\Delta\mathbf{w}$ were [0.06, 0.95, 0.44] degrees, all below 1.0 degrees. Similarly, for the TL-shaped object, the deviations in $\Delta\mathbf{p}$ and $\Delta\mathbf{w}$ were [1.86, 1.50, 0.36] mm and [0.47, 0.38, 0.85] degrees, respectively, remaining below 2 mm and 0.9 degrees. These results suggest that while the overall estimation errors appear large, their internal consistency is high, indicating that the system’s inaccuracies stem largely from systematic biases. Likely sources include errors from the optical tracking system itself (due to the 3D-printed marker mounts for calibration and robot frame location), installation errors in the dual-arm robot, and absolute control errors of the robot. In addition, when comparing the deviations between the L-shaped and TL-shaped objects, we found that the inference results for the L-shaped object were better than those for the TL-shaped object. This outcome differs from our theoretical expectations. Theoretically, even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations. We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames, which place higher demands on the assembly precision of the dual-arm robot. The larger deviations reveal that the robot setup has greater misalignment in rotational components.
@@ VII Conclusions and Future Work
The current method assumes planar contacts between flat fingertips and flat object surfaces. When the object surface is curved, the number of degrees of freedom constrained by a grasp will vary, and more grasps may be required to constrain uncertainties. Investigating the minimal and sufficient set of grasps under such conditions is an interesting future direction. Additionally, exploring learning-based methods to predict grasps and motion will be helpful to improve efficiency.

### v4

@@ Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concept that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three approximately orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three approximately orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that it had good repeatability. The deviation levels of the experimental trials were on the same order of magnitude as those of an optical tracking system, demonstrating strong relative inference performance.
@@ I Introduction
Figure 1: (a) Precisely grasping an object is difficult. (a.i) Before grasping. (a.ii) After grasping. (a.iii) Comparison. (b) Flat finger pads of a parallel gripper help constrain uncertainty to a plane. (c) Three aprpoximately orthogonal parallel grasps constrain three planes, leading to an exact pose.
This paper proposes a bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to reduce uncertainties. It does not require external fixtures or additional sensors and is more flexible for diverse robotic manipulation scenarios. The fundamental concept of the proposed method is illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state Note 1 Note 1 Note 1This unique pose constraint holds under the following assumptions: i) The object is sufficiently rigid such that the grasping force is transmitted and sensed through the force sensors, rather than causing deformation. ii) The contact between the gripper fingers and the object is assumed to be flat surface contact with sufficient friction to prevent slippage. iii) The maximum uncertainty in the object pose is within the gripper’s opening range, ensuring valid contact during grasping.. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting approximately orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors at the robots’ wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three approximately orthogonal axes, thus achieving an accurate object pose estimation.
@@ II-A Categories of Regrasp Planning Methods
Continuous regrasp methods, on the other hand, aim to adjust object poses through controlled rolling or sliding motions while maintaining contact. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds.
The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it falls under the first type of regrasp planning. It can be seen as a concrete realization of a flipped funnel [17], where the object, through interaction, progressively guides the three structured grasping poses, allowing its uncertain state to be estimated toward a determined pose.
@@ III Regrasp Planning Using Orthogonal Triplets
Figure 2: (a) Antipodal grasp planning. (b) Candidate grasp poses planned using the antipodal method. (c) Separating the candidate grasp poses into three orthogonal groups considering their opening and closing directions. Each color in the subfigure represents one group.
Subsequently, based on the results of the grasp planning, we separate the grasps into groups. Since our goal is to identify three approximately orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 2 Note 2 Note 2In this context, the direction of $\boldsymbol{v}_{i}$ is not considered.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …. The candidate grasp poses in Fig. 2(b) are classified into three strictly orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue.
After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the approximately orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triplet combination.
@@
It is important to note that strict 90-degree orthogonality is not required to determine the object’s pose. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triplet combinations. The triplets may have non-zero scores, However, as long as the grasp direction vectors $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ form a non-singular matrix when arranged as columns, the pose can be estimated. This point will be clearer in Section V when we introduce the pose estimation algorithm. Here, the reason we sort triplet grasp combinations by their score is not for enforcing strict orthogonality, but rather to account for potential changes of grasp directions caused by subsequent admittance control. Prioritizing grasp triplets with higher orthogonality helps minimize the impact of grasp direction deviations. We will discuss this aspect in more detail in the admittance control section.
After identifying the approximately orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. We use an incremental method to explore combinations of grasp triplets from the orthogonal groups and output a successful motion once found. Each triplet comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triplets and output successful motion once found. The detailed algorithmic flow is shown in Fig. 3.
@@
It should be noted that our grasping and planning methods are probabilistic, and their success rates inherently depend on the granularity, number, and diversity of the grasp samples. At the same time, increasing the number of grasps inevitably impacts planning time. Identifying the optimal trade-off between these factors is an important challenge that could be addressed using predictive approaches. If a large amount of data can be accumulated using the proposed method in the future, we are interested in exploring probabilistic modeling methods to learn grasp triplets and thus improve planning efficiency. However, this is beyond the current scope and will not be elaborated.
@@ VI Experiments and Analysis
Table I shows the results. The data section has two main column groups reporting the positional differences and rotational differences. Within each group, there are three subcolumns. The “second grasp” columns show the positional and rotational differences of the $g_{2}$ grasp before and after admittance control. The “third grasp” columns present the corresponding differences for the $g_{3}$ grasp. The “object” columns report the differences between the object pose estimated by the proposed method and the pose measured by the optical tracking system. The “first grasp” uses position control and is thus excluded.
@@
Position Differences ($\Delta\mathbf{p}$, mm) Rotation Vector Differences ($\Delta\mathbf{w}$, degrees)
second grasp third grasp object second grasp third grasp object
@@
Upon closer examination, we found that these errors are primarily systematic. For the L-shaped object, the standard deviations in $\Delta\mathbf{p}$ across tasks were [0.25, 0.15, 0.25] mm, all below 0.3 mm, and the deviations in $\Delta\mathbf{w}$ were [0.06, 0.95, 0.44] degrees, all below 1.0 degrees. Similarly, for the TL-shaped object, the deviations in $\Delta\mathbf{p}$ and $\Delta\mathbf{w}$ were [1.86, 1.50, 0.36] mm and [0.47, 0.38, 0.85] degrees, respectively, remaining below 2 mm and 0.9 degrees. These results suggest that while the overall estimation errors appear large, their internal consistency is high, indicating that the system’s inaccuracies stem largely from systematic biases. Likely sources include calibration errors in the optical tracking system, installation errors in the two robot arms and the mounting aluminum frame, and absolute control errors of the robots. In addition, when comparing the deviations between the L-shaped and TL-shaped objects, we found that the inference results for the L-shaped object were better than those for the TL-shaped object. This outcome differs from our theoretical expectations. We can conclude from the equations in Section V that even if the three grasps are not perfectly orthogonal, the final inference accuracy should not degrade, and the L- and TL-shaped objects should exhibit similar deviations. We believe that the observed difference arises because the TL-shaped object involves rotations around axes unaligned with body frames. The large difference indicates that the robot setup has greater rotational misalignment. This raises the question of compensating for system bias to improve absolute accuracy. While theoretically feasible, such correction would require complex estimation of both positional and rotational errors, which we leave for future work.
@@ VII Conclusions and Future Work
In our experiments, we used a dual-arm setup built from two single-arm robots mounted on an aluminum frame, which inevitably introduced systematic errors and led to relatively large absolute errors. As a next step, we plan to validate the method in a dynamic simulation environment, where no absolute error is present, and use a dual-arm robot with commercially-calibrated body integration for practical deployment.